OpenStudy (anonymous):

for tan = y/x what does it mean that x does not equal 0 and there is a restriction at 90 and 270?

OpenStudy (anonymous):

You can't divide by zero. What would computing the tangent of 90 mean (think of a right angled triangle) Same problem for 270 (= -90 degrees).

OpenStudy (anonymous):

hmmm storta confused..

OpenStudy (anonymous):

OK What's your calculator say for tan 89, tan 89.9, tan 89.99 etc

OpenStudy (anonymous):

57.28

OpenStudy (turingtest):

...or how about what is the definition of \(\tan x\)

OpenStudy (turingtest):

?

OpenStudy (turingtest):

@HorribleAtMath Why run away? how does that help you

OpenStudy (anonymous):

One definition is opposite/adjacent sides of a right hand triangle - draw a sketch.

OpenStudy (anonymous):

i didn't run away. geese I'm on a different question. tan is y/x

OpenStudy (turingtest):

I was going to say\[\tan x=\frac{\sin x}{\cos x}\]right?

OpenStudy (anonymous):

except when cosx = 0

OpenStudy (turingtest):

...that's what I'm getting at...

OpenStudy (anonymous):

so like if i have sin 3x = 1/2

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

You say what angle has a sine of a half, and go from there.

OpenStudy (anonymous):

for sin3x = 1/2 why is the answer pi/18 and 5pi/18 ?

OpenStudy (anonymous):

You are using radians to measure angles.

OpenStudy (anonymous):

can you just answer the question

OpenStudy (anonymous):

if the sine of something = 1/2, something must be 30 degrees (or 120 degrees) or add a multiple of 360 degrees...

OpenStudy (anonymous):

Correction; read 150, not 120.